\begin{thebibliography}{}

%\bibitem {alekseev} Alekseev, M. V.: Quasirandom diffeomorphisms I, Math. USSR Sbornik \textbf{5(73)} (1968)

%\bibitem {barreira} Barreira, L.: A non additive thermodynamic formalism and applications to dimension theory of hyperbolic dynamical systems. Erg. Theory \& Dyn. Sys. \textbf{16} (1996) 871-927 

%\bibitem {barreira.gelfert} Barreira, L. - Gelfert, K.: Dimension estimates in dynamical systems: a survey of recent results, Erg. Th. \& Dyn. Sys. \textbf{31} (2011), 641-671

%\bibitem {barreira.2010} Barreira, L.: Thermodynamic Formalism and Applications to Dimension Theory, Progress in Mathematics Vol. 294, Birkh\"auser, Springer, Basel (2011)

\bibitem {barreira.pesin} Barreira, L. - Pesin, Ya.: Non Uniform Hyperbolicity, Encyclopedia of Mathematics and its Applications, Cambridge University Press, (2007)

\bibitem {bonatti.baladi.schmitt} Bonatti, C. - Baladi, V. - Schmitt, C.: Abnormal escape rates from nonuniformly hyperbolic sets, Erg. Th. \& Dyn Sys 19(05) (1999) 1111-1125

%\bibitem {bowen.1977}  Bowen, R. : On Axiom A diffeomorphisms. Conference Board Math. Sci. \textbf{33}, American Mathematical Society, Providence (1977)

\bibitem {bowen} Bowen, R.: Equilibrium States and the Ergodic Theory of Anosov Diffeomorphisms, Lecture Notes in Math. 409, Springer Verlag, New York (1975)

\bibitem {cao.luzzatto.rios} Cao, Y. - Luzzatto, S. - Rios, I.: Some non-hyperbolic systems with strictly non-zero Lyapunov exponents for all invariant measures: horseshoes with internal tangencies,  Disc. \& Continuous Dyn. Sys. \textbf{15(1)} (2006)

\bibitem {chernov}  Chernov, N.; Markarian, R.; Troubetzkoy, S. Invariant measures for Anosov maps with small holes, Erg. Th. \& Dyn Sys  20(4) (2000), 1007-1044.

\bibitem {demers.young} Demers, M. - Young, L.S.: Escape rates and conditionally invariant measures, Nonlinearity \textbf{19} (2006) 377-397.

\bibitem {demers} Demers, M.: Escape rates and physical measures for the infinite horizon Lorentz gas with holes, Dynamical Systems: An International Journal \textbf{28(3)} 2013, 393-422.

%\bibitem {bruin.todd} Bruin, H. - Todd, M.: Equilibrium States for Interval Maps: Potentials with $\sup\phi - \inf\phi < h_{top}(f)$, Commun. Math. Phys. \textbf{283} (2008) pp. 579-611

%\bibitem {dolgopyat.pesin} Dolgopyat, D. - Pesin, Ya.: Every compact manifold carries a completely hyperbolic diffeomorphism, Erg. Th. \& Dyn Sys, 22 (2002), 409-435.

%\bibitem {cao.feng.huang} Cao, Y.-L.-Feng, De-Jun, Huang, W.: The thermodynamic formalism for subadditive potentials, Disc. \& Cont. Dyn. Sys. \textbf{20(3)} (2008) 639-657

\bibitem {eckmann.ruelle.1985} Eckmann, J.P. - Ruelle, D.: Ergodic theory of chaos and strange attactors, Review of Modern Physics, \textbf{57(3)} (1985), 617-656

\bibitem {gonzalo} Contreras-Barandari\'an, G.: Rate of Escape of Some Chaotic Julia Sets, Commun. Math. Phys. 133 (1990) 197-215  

%\bibitem {falconer} Falconer, J. : A subadditive thermodynamic formalism for mixing repellers, J. Phys. A: Math. Gen. \textbf{21} (1988) L737-L742

%\bibitem {gelfert.wolf} Gelfert, K. - Wolf, C.  On the distribution of periodic orbits. Disc. \& Cont. Dyn. Sys., Series A \textbf{26(3)} (2010) 949-966. 

%\bibitem {gelfert.2009} Gelfert, K.: Equality of pressures for diffeomorphisms preserving hyperbolic measures, Math. Z. \textbf{261(4)} (2009) 711-724 

%\bibitem {gelfert.2010} Gelfert, K.: Repellers for non-uniformly expanding maps with singular or critical points, Bulletin of the Brazilian Mathematical Society, New Series 41(2) (2010) 237-257

%\bibitem {guckenheimer.holmes} Guckenheimer, J. - Holmes, J.: Nonlinear Oscillations, Dynamical Systems, and Bifurcations of Vector Fields (Applied Mathematical Sciences) 42, Springer-Verlag, New York 1983.

%\bibitem {gurevic.1998} B. M. Gurevich B.M. - Savchenko, S.V. : Thermodynamic formalism for countable symbolic Markov chains, Russian Math. Surveys \textbf{53:2} (1998) 245-344

%\bibitem {hofbauer.keller} Hofbauer, F. - Keller, G.: Equilibrium states for piecewise monotonic transformations, Erg. Theory \& Dyn. Sys. \textbf{2(01)} (1982) pp. 23-43.

%\bibitem {katok} Katok, A. : Lyapunov exponents, entropy and periodic orbits of diffeomorphisms, Publ. Math. IHES \textbf{51} (1984) 137-173 

%\bibitem {katok.hasselblat} Katok, A. - Hasselblat, B.: Introduction to the Modern Theory of Dynamical Systems, by Katok, A. and Hasselblat, B., Encyclopedia of Math. and its Applications. Vol. 54. Cambridge University Press, 1995 

%\bibitem {katok.mendoza} Katok, A. - Mendoza, L.: Dynamical Systems with Nonuniformly Hyperbolic Behavior. Supplement to Introduction to the Modern Theory of Dynamical Systems, by Katok, A. and Hasselblat, B., Encyclopedia of Math. and its Applications. Vol. 54. Cambridge University Press, 1995 

%\bibitem {kingman} Kingman, J.F.C.: The subadditive ergodic theory, Ann. of Probability, \textbf{1(6)} (1973) 883-909

%\bibitem {liang.liu.sun} Liang, Ch.-Liu, G.-Sun, W.: Approximation property on invariant measure and Oseledets splitting in nonuniformly hyperbolic systems, Trans. of the Amer. Math. Soc. \textbf{361(3)} (2009) 1543-1579
  
\bibitem {lopes.markarian} Lopes, A. - Markarian, R.: Open billiards: invariant and conditionally invariant probabilities on Cantor sets, Siam J. Appl. Math. 56 (1996), 651-680.

\bibitem {luzzatto.sanchez} Luzzatto, S. - S\'anchez-Salas, F.: Uniform hyperbolic approximations of measures with non zero Lyapunov exponents, Proc. of the Amer. Math. Soc. \textbf{141(9)} (2013) 3157-3169

%\bibitem {manhe} Ma\~n\'e, R. : Ergodic Theory of Differentiable Dynamics, Series of Modern Surveys in Mathematics, Volume 8, Ergebnisse Der Mathematik Und Ihrer Grenzgebiete. 3. Folge a, Springer Verlag, London (1987)

%\bibitem {mendoza.1985} Mendoza, L.: The entropy of $C^2$ surface diffeomorphisms in terms of Hausdorff dimension and a Lyapunov exponent, Erg. Theory \& Dyn. Sys. \textbf{5} (1985) 273-283

%\bibitem {mendoza.1988} Mendoza, L. : Ergodic attractors for diffeomorphisms of surfaces, J. London Math. Soc. \textbf{s2-37(2)} (1988) 362-374 

\bibitem {mora.viana} Mora, L. - Viana, M.: Abundance of strange attractors, Acta Math., \textbf{171} (1993), 1-71.

%\bibitem {misiurewicz} Misiurewicz, M.: A short proof of the variational principle for $\enteros_+^{\naturales}$ action on a compact space, Asterisque \textbf{40} (1976) 147-157

%\bibitem {newhouse} Newhouse, S. : Continuity properties of the entropy, Ann. of Math. \textbf{129(1)} (1989) 215-235 

%\bibitem {pesin.pitskel} Ya. B. Pesin and B. S. Pitskel': Topological pressure and the variational principle for non-compact sets, Functional Anal. Appl. \textbf{18} (1984), 5063-

%\bibitem {oliveira.2003} Oliveira, K.: Equilibrium states for nonuniformly expanding maps, Erg. Theory \& Dyn. Sys. \textbf{23(06)} (2003) pp. 1891-1905.

%\bibitem {palis.yoccoz} Palis, J. - Yoccoz, J.C.: Non-uniformly hyperbolic horseshoes arising from bifurcations of Poincaré heteroclinic cycles, Publications Math\'ematiques I.H.E.S., November 2009, \textbf{110(1)}, pp 1-217 

\bibitem {pesin.2004} Pesin, Ya.: Thermodynamics of systems with nonzero exponents more questions than answers, Dynamics of Complex Systems (kurims.kyoto-u.ac.jp) 2004, vol. 1404, p. 91-97.

\bibitem {pesin.2010} Climenhaga, V. - Pesin, Ya.: Open problems in the theory of nonuniform hyperbolicity, Disc. \& Continuous Dyn. Sys. \textbf{27(2)} (2010). 

%\bibitem {ruelle.1973} Ruelle, D: Statistical mechanics on a compact set with $\enteros^n$ action satisfying expansiveness and specification, Trans. of Amer. Math. Soc. \textbf{187} (1973), 237-251

%\bibitem {ruelle.1979} Ruelle, D.: The ergodic theory of differentiable dynamical systems, Publ. Math. IHES \textbf{50(1)} (1979)

\bibitem {ruelle.1996} Ruelle, D.: Positivity of entropy production in nonequilibrium statistical mechanics, Journal of Statistical Physics, \textbf{85(1,2)} (1996)

%\bibitem {pesin} Pesin, Ya.: Dimension theory in Dynamical Systems, Chicago Lecture Notes in Mathematics, Contemporary Views and Applications, The University of Chicago Press, Chicago, 1997

%\bibitem {sanchez-salas.1} Fernando J. S\'anchez-Salas: Ergodic attractors as limits of hyperbolic horseshoes, Erg. Theory \& Dyn. Sys. \textbf{22} (2002) 571-589

%\bibitem {sanchez-salas.2} Fernando J. S\'anchez-Salas: Dimension of Markov towers for non uniformly expanding one dimensional system, Disc. \& Continuous Dyn. Sys. \textbf{9(6)} (2003)

\bibitem {sanchez.salas.2015} Fernando J. S\'anchez-Salas: On the approximation of dynamical indicators in systems with nonuniformly hyperbolic behavior, arXiv:1505.02473 (2015).

%\bibitem {tucker} Tucker, W.: A Rigorous ODE Solver and Smale’s 14th Problem, Found. Comput. Math. (2002) \textbf{2}, pp. 53-117.

%\bibitem {varandas.viana.2009} Varandas, P. - Viana, M.: Existence, uniqueness and stability of equilibrium states for nonuniformly expanding maps, Ann. de L'Institute Henri Poincaré \textbf{27} (2010) pp. 555-593

%\bibitem {viana} Viana, M.: Strange attractors in higher dimensions, Bol. Soc. Brasil. Mat. (New Series), \textbf{24} (1993), 13-62.

%\bibitem {walters} Walters, P.: A variational principle for the pressure of continuous transformations, Amer. Journal of Math., \textbf{97(4)} (1975) 937-971

%\bibitem {young.1981} Young, L. -S.: On the prevalence of horseshoes, Trans. Amer. Math. Soc. \textbf{263} (1981), 75-88

%\bibitem {young.1982} Young, L.-S.: Dimension, entropy and Lyapunov exponents, Erg. Theory \& Dyn. Sys. \textbf{2} (1982), 109–124

\bibitem {young.1990} Young, L. S.: Some large deviations results for dynamical systems, Trans. Amer. Math. Soc. \textbf{318} (1990), 525 - 543

\bibitem {young.1998} Young, L. S.: 

%\bibitem {wang.sun} Wang, Z. - Sun, W.: Lyapunov exponents of hyperbolic measures and hyperbolic periodic orbits, Trans. of the Amer. Math. Soc. \textbf{362(8)} (2010) 4267-4282



\end{thebibliography}

\documentclass[10pt]{amsart}

\usepackage[a4paper]{geometry}

\usepackage{graphicx}
\usepackage{amsmath,amsthm, amsfonts, amssymb}

\newtheorem*{theorem*}{Theorem}
\newtheorem{lemma}{Lemma}[section]
\newtheorem{proposition}[lemma]{Proposition}

\theoremstyle{definition}
\newtheorem{definition}[lemma]{Definition}
\theoremstyle{remark}

\newcommand{\dist}{\mbox{dist}}

\newcommand{\Vol}{\mbox{Vol}\, }

\title[Regular nonuniformly hyperbolic systems]{A note on Eckmann-Ruelle's conjecture}

\author{ 
Fernando Jos\'e S\'anchez-Salas}
\address{
Departamento de Matem\'aticas, Facultad Experimental de Ciencias, Universidad del Zulia, Avenida Universidad, Edificio Grano de Oro, Maracaibo, Venezuela
}
\email{fjss@fec.luz.edu.ve}

\date{April 8, 2015}

\subjclass[2010]{37D25, 37D35}

\keywords{Nonuniformly hyperbolic systems, rates of escape, hyperbolic product structure}

\thanks{ This work was partially supported by the Associateship Programme of ICTP}

\begin{document}

\begin{abstract}
We introduce a class of $C^{1+\alpha}$ isolated nonuniformly hyperbolic sets $\Lambda$ for which $\sup_{\mu \in {\mathcal M}_f}\{h(\mu) - \chi^+(\mu)\}$ 
equals the rate of escape from $\Lambda$, where $\chi^+(\mu)$ is the average of the sum of positive Lyapunov exponents counted with their multiplicity.
\end{abstract}

\maketitle

\section{Introduction}\label{sec:introduction}

Let $\Lambda$ be a compact $f$-invariant isolated subset, that is, there exists a neighborhood $U$ containing $\Lambda$, such that 
$\Lambda = \bigcap_{-\infty}^{+\infty}f^n(U)$. The {\em rate of escape of $\Lambda$ from $U$} is defined as
\begin{equation}\label{rate.escape.1}
\rho(U) := \limsup_{m \to +\infty}\dfrac{1}{m}\log(\Vol(U_m)),
\end{equation}
where $$U_m = \{\,x \in U \ :\  f^k(x) \in U \ \text{for} \ k = 0, \cdots , m\,\}$$ is the set of points in $U$ 
which stay in $U$ up-to its first $m$ iterates.

It is well known from the thermodynamics of $C^{1+\alpha}$ Axiom A systems that the rate of escape from a suitable small neighborhood $U$ of a topologically transitive 
isolated uniformly hyperbolic set $\Omega$ equals the topological pressure of the unstable potential $\phi^u = -\log(Jac(Df|E^u)$:
$$
P(f | \Omega,\phi^u) = \rho(U).
$$
See \cite{bowen}. By the variational principle for topological pressure 
$$
P(f | \Omega,\phi^u) = \sup\limits_{\mu \in {\mathcal M}_f}\{h(\mu) - \chi^+(\mu)\} = \rho(U),
$$
where
$$
\chi^+(\mu) = \int\sum_{\chi_i(x) > 0}\chi_i(x)\dim(E_i(x))d\mu(x)
$$ 
the average of the sum of positive Lyapunov exponents counted with their multiplicity. Also by the thermodynamics of $C^{1+\alpha}$ Axiom A systems there always exists a unique ergodic 
Borel probability $\mu = \mu_{\phi^u}$ which is an equilibrium state for $\phi^u$ and therefore
$$
h(\mu) - \chi^+(\mu) = \rho(U).
$$

Numerical evidence and heuristic arguments suggested that this property holds for general isolated compact $f$-invariant subsets (see \cite{bonatti.baladi.schmitt}) leading 
Eckmann and Ruelle to raise the following
\ 
\\
\\
{\bf Conjecture} \ {\em Let $\Lambda \subset M$ be a locally maximal compact $f$-invariant set of a smooth diffeomorphism of a compact Riemannian manifold and suppose that 
there exists an ergodic measure $\mu$ such that 
\begin{equation}\label{generalized.SRB}
h(\mu) - \chi^+(\mu) = \sup\limits_{\nu \in {\mathcal M}_f}\{h(\nu) - \chi^+(\nu)\}.
\end{equation}
Then $h(\mu) - \chi^+(\mu) = \rho(U)$ is the rate of escape of a sufficiently small isolating neighborhood $U$ of $\Lambda$,   
}

See \cite{eckmann.ruelle.1985}. An extreme $\mu$ of the variational equation (\ref{generalized.SRB}) is known as a generalized Sinai-Ruelle-Bowen (SRB) 
measure. See \cite{ruelle.1996}. It is not known how generally the conjecture holds beyond systems satisfying Axiom A. 
Eckmann-Ruelle conjecture has been proved for uniformly partially hyperbolic diffeomorphisms \cite{young.1990}, Julia sets of rational maps of the 
Riemann sphere \cite{gonzalo} and certain billiars and so called open systems in \cite{chernov}, \cite{demers.young}, \cite{demers}, 
\cite{lopes.markarian}. We refer the reader to \cite{bonatti.baladi.schmitt} for a review of the problem.

In \cite{bonatti.baladi.schmitt} Baladi, Bonatti and Schmitt constructed a nonuniformly hyperbolic counterexamples to 
Eckmann-Ruelle's conjecture, that is: compact invariant nonuniformly hyperbolic locally maximal set $\Lambda$ with an isolating neighborhood $U$
and an ergodic (unique) non atomic hyperbolic measure $\mu$ such that $\Lambda = supp \ \mu$,
$$
\rho(U) = 0 \quad\text{and}\quad h(\mu) - \chi^+(\mu) = \sup\limits_{\nu \in {\mathcal M}_f}\left\{\,h(\nu) - \chi^+(\nu)\,\right\} < 0.
$$
The example is made from a plug of the eye-like Bowen example into an uniformly hyperbolic repellor $\Lambda_0$ after blowing up a 
suitable fixed point $x_0 \in \Lambda_0$. Then they insert into the blow up a configuration $\Lambda_1$ of finitely many hyperbolic fixed points with 
saddle connections including a copy of a well known Bowen's example made of two hyperbolic fixed points with saddle connections in a eye-like form. 
However the invariant set formed by the two separatrices in Bowen's example do not support an invariant measure so, in principle, it does not gives a 
counterexample for the Eckmann-Ruelle conjecture. Then they handle the plugin in such way that $\mu_0$ the unique generalized SRB measure for 
$\Lambda_0$ extends to a non atomic ergodic measure $\mu$ with support $\Lambda = \Lambda_0 \cup \Lambda_1$ which is also an SRB measure for $\Lambda$ 
and $h(\mu) - \chi^+(\mu) < 0$.

We will prove that, except for this type of examples, Eckmann-Ruelle's conjecture holds for {\em regular nonuniformly hyperbolic sets}.

\begin{definition}\label{definition.regular.nonuniformly.hyperbolic}
Let $\Lambda$ be a compact $f$-invariant nonuniformly hyperbolic set of a $C^{1+\alpha}$ diffeomorphism of a compact manifold. We say that $\Lambda$ is 
a \emph{regular nonuniformly hyperbolic set} if $\Lambda$ is isolated and there exists an increasing sequence of basic sets $\Lambda_n \subset \Lambda$ such that:
\begin{enumerate}
\item $\Lambda = \overline{\bigcup_n\Lambda_{n}}$ and
\item $U = \bigcup_nU_n$ is an isolating open neighborhood of $\Lambda$, where $U_n \subset U_{n+1}$ is an increasing sequence of open isolating neighborhoods of 
$\Lambda_n$.
\end{enumerate}
\end{definition}

We call such a sequence a \emph{regular exhaustion of $\Lambda$}.
\ 
\\
\\
{\bf Theorem A} \ {\em Let $\Lambda \subset M$ be a regular nonuniformly hyperbolic set of a $C^{1+\alpha}$ diffeomorphism. Then,
\begin{equation}\label{eckmann.ruelle.1}
\sup\limits_{\mu \in {\mathcal M}_f}\{h(\mu) - \chi^+(\mu)\} = \rho(U).
\end{equation}
}
\ 
\\
All known nonuniformly hyperbolic sets which are known to verify the Eckmann-Ruelle conjecture are regular. See  \cite{chernov}, \cite{demers.young}, \cite{demers}, 
\cite{lopes.markarian}. We conclude from Theorem A that if $\Lambda$ is a regular nonuniformly hyperbolic isolated set and it supports a nonatomic 
generalized SRB measure then it satisfies the Eckmann-Ruelle conjecture. Theorem A follows from definition \ref{definition.regular.nonuniformly.hyperbolic} 
and the following
\ 
\\
\\
{\bf Theorem B} \ {\em Let $\Lambda$ be a regular nonuniformly hyperbolic set of a $C^{1+\alpha}$ diffeomorphism of a compact manifold. Then,
\begin{equation}\label{eckmann.ruelle.2}
\sup\limits_{\mu \in {\mathcal M}_f(\Lambda)}\{h(\mu) - \chi^+(\mu)\} = \sup\limits_{\Omega \in {\mathcal H}}\rho(U_{\Omega}),
\end{equation}
where ${\mathcal M}_f(\Lambda)$ is the set of $f$-invariant Borel probabilities in $\Lambda$, $\mathcal{H}$ is the family of basic sets $\Omega \subset \Lambda$ and $\rho(U_{\Omega})$ denote the rate of escape from a small isolating neighborhood 
$U_{\Omega}$ of $\Omega \in {\mathcal H}$.
}
\ 
\\
\begin{proof}[Proof of Theorem A]
Let $\Lambda_n \subset \Lambda$ be a regular exhaustion and $U_n$ the corresponding isolating neighborhoods. We shall prove that:
\begin{equation}\label{theoremA.1}
 \rho(U) = \sup_{n > 0}\rho(U_n).
\end{equation}
First we see that (\ref{theoremA.1}) proves Theorem A. By the Theorem B
$$
\rho(U) = \sup_{n > 0}\rho(U_n) \leq \sup\limits_{\Omega \in {\mathcal H}}\rho(U_{\Omega}) = \sup\limits_{\mu \in {\mathcal M}_f(\Lambda)}\{h(\mu) - \chi^+(\mu)\}.
$$
Therefore, $\rho(U) \leq \sup_{\mu \in {\mathcal M}_f}\{h(\mu) - \chi^+(\mu)\}$. The inequality 
$\rho(U) \geq \sup_{\mu \in {\mathcal M}_f}\{h(\mu) - \chi^+(\mu)\}$ follows from L.S. Young's \cite{young.1990} results.

To prove (\ref{theoremA.1}) we first notice that $U^{(m)} = \bigcup_nU^{(m)}_n$ implies $\Vol(U^{(m)}) = \sup_{n > 0}\Vol{U^{(m)}_n}$ and then, as $\Vol{U^{(m)}_n} \leq \Vol{U^{(m)}_{n+1}}$ 
for every $n > 0$,
\begin{eqnarray*}
 \limsup_{m \to +\infty}\dfrac{1}{m}\log(\Vol{U^{(m)}}) & \geq &  \limsup_{m \to +\infty}\dfrac{1}{m}\log\left(\sup_{n > 0}\Vol{U^{(m)}_n}\right)\\
                                                        &  =   & \limsup_{m \to +\infty}\dfrac{1}{m}\sup_{n > 0}\log(\Vol{U^{(m)}_n)}\\
                                                        &  =   & \inf_{m > 0}\sup_{k \geq m}\dfrac{1}{k}\sup_{n > 0}\log(\Vol{U^{(k)}_n})\\
                                                        &  =   & \inf_{m > 0}\sup_{n > 0}\sup_{k \geq m}\dfrac{1}{k}\log(\Vol{U^{(k)}_n}),\\
\end{eqnarray*}
Let us denote
$$
V^{(m)}_n = \sup_{k \geq m}\dfrac{1}{k}\log(\Vol{U^{(k)}_n}).
$$
Then, $\rho(U_n) = \inf_{m > 0}V^{(m)}_n$. Moreover, let 
$$
V^{(m)} = \sup_{k \geq m}\dfrac{1}{k}\log(\Vol{U^{(k)}}).  
$$
Therefore,
$$
V^{(m)} = \sup_{n > 0}V^{(m)}_n \geq \sup_{n > 0}\inf_{m > 0}V^{(m)}_n = \sup_{n > 0}\rho(U_n), \quad\forall \ m > 0.
$$
Hence 
$$
\rho(U) = \inf_{m > 0}V^{(m)} \geq \sup_{n > 0}\rho(U_n).
$$
Now let $\epsilon > 0$ small and choose, for every $n > 0$ an integer $m_n > 0$ such that
\begin{equation}\label{theoremA.2}
 V^{(m_n)}_n < \inf_{m > 0}V^{(m)}_n + \epsilon \quad\forall \ n > 0.
\end{equation}
Then, for every small $\epsilon > 0$.
\begin{eqnarray*}
 \rho(U) = \inf_{m > 0}V^{(m)} & = & \inf_{m > 0}\sup_{n > 0}V^{(m)}_n \\
                               & \leq & \sup_{n > 0}V^{(m_n)}_n \\
                               & \leq & \sup_{n > 0}\inf_{m > 0}V^{(m)}_n + \epsilon\\
                               &  =   & \sup_{n > 0}\rho(U_n) + \epsilon,
\end{eqnarray*}
and we conclude that
$$
\rho(U) \leq \sup_{n > 0}\rho(U_n).
$$
\end{proof}

To prove of Theorem B we use the following

\begin{proposition}\label{prop.1}
$\Lambda$ be a $C^{1+\alpha}$ nonuniformly hyperbolic compact $f$-invariant set. Suppose in addition that there exists an increasing sequence of basic sets 
$\Lambda_n \subset \Lambda$ such that $\Lambda = \overline{\bigcup_n\Lambda_n}$. Then, for every continuous $\phi$
\begin{equation}\label{main.1}
P(f|\Lambda, \phi) = \sup_{\Omega \in \mathcal{H}}P(f|\Omega,\phi)
\end{equation}
where $ \mathcal{H}$ is the family of basic sets $\Omega \subset \Lambda$. 
\end{proposition}

We proved in \cite{sanchez.salas.2015} that (\ref{main.1}) holds in general for every nonuniformly hyperbolic systems and an special class of 
so called \emph{hyperbolic potentials}. Actually we show that there are nonuniformly hyperbolic systems for which (\ref{main.1}) does not hold for every 
continuous potential. This is the case if the dynamics has isolated hyperbolic orbits as is the case, for example for a diffeomorphism of the sphere having a horseshoe with 
internal tangencies and a repeller point at north pole, as we show in \cite{sanchez.salas.2015}. So proposition \ref{prop.1} is interesting since it 
gives a sufficient condition for the variational equation (\ref{main.1}) to hold for \emph{every continuous function}.  

If $\Lambda$ is the support of a nonatomic hyperbolic measure then there exists a sequence of uniformly hyperbolic sets $\Lambda_n \subset \Lambda$ 
such that $\Lambda = \overline{\bigcup_n\Lambda_n}$. This follows from \cite{luzzatto.sanchez}. Notice however that this is not a sufficient
condition for $\Lambda$ to has a regular exhaustion, as Baladi et al example shows.

Proposition \ref{prop.1} follows from the continuity of topological pressure $P(f|\Lambda,\phi)$ with respect to $\Lambda$ and the existence of an 
exhausting sequence $\Lambda_n$.

\begin{definition}
Let $X$ be a compact metric space. We define the Hausdorff distance of compact subsets $A,B \subset X$ as:
$$
\dist_H(A,B) = \inf\{\epsilon > 0: A \subset B_{\epsilon}, \ B \subset A_{\epsilon}\}
$$
where $A_{\epsilon} := \{x \in X: \dist(x,A) < \epsilon\}$ and $\dist(x,A) = \inf_{y \in A}d(x,y)$.
\end{definition}

Let $f : X \to X$ be a continuous selfmap of a compact metric space. The family $\mathcal{I}$ of compact $f$-invariant subsets is 
a compact metric space with the Hausdorff distance.

\begin{lemma}\label{lemma.1}
Let $\phi$ be continuous and $f : X \to X$ be a continuous selfmap of a compact metric space. Then $\Omega  \mapsto P(f|\Omega,\phi)$ is continuous when 
$\Omega$ varies on the family of compact $f$-invariant subsets with $P(f|\Omega,\phi) < +\infty$.
\end{lemma}
\begin{proof}
$$
P(f|\Omega,\phi,\epsilon) = \limsup_{n \to +\infty}\dfrac{\log{P_n(f,\phi,\epsilon)}}{n}
$$
where
$$
P_n(f,\phi,\epsilon) = \sup_E\sum_{x \in E}e^{S_n\phi(x)},
$$
infimun taken over the familiy of $(\epsilon,n)$-separated sets and $S_n\phi(x) = \sum_{k=0}^{n-1}$. Then, given $\delta > 0$, let 
$0 < \epsilon < \delta$ such that if $d(x,y) < \epsilon$ then $|\phi(x) - \phi(y)| < \delta/2$ and $n > 0$ such that
$$
\exp{n(P(f|\Omega,\phi,\epsilon/2) - \delta/2)} < P_n(f,\phi,\epsilon/2) <  \exp{n(P(f|\Omega,\phi,\epsilon/2) + \delta/2)}.
$$
Let $E \subset \Omega$ be a maximal $(\epsilon/2,n)$-separated set. Then, $\Omega \subset \bigcup_{x \in E}B(x,n,\epsilon/2)$. Choose $\eta > 0$ 
sufficiently small such that $d(x,y) < \eta$ implies $d(f^k(x),f^k(y)) < \epsilon/2$ for $k = 0, \cdots , n-1$ and such that 
$\dist_{H}(\Omega,\Omega') < \eta$ implies $\Omega' \subset \bigcup_{x \in E}B(x,n,\epsilon/2)$. For every $x \in E$ we choose $x' \in \Omega'$ such that
$d(x,x') < \eta$ and call $F$ such set. Then $\Omega' \subset \bigcup_{x' \in F}B(x',n,\epsilon)$ and
$$
e^{-n\delta/2}\sum_{x \in E}e^{S_n\phi(x)} < \sum_{x' \in F}e^{S_n\phi(x')} < e^{n\delta/2}\sum_{x \in E}e^{S_n\phi(x)}.
$$
Then,
$$
Q_n(f|\Omega',\phi,\epsilon) = \inf_F\sum_{x \in F}e^{S_n\phi(x)} < \exp{n(P(f|\Omega,\phi,\epsilon/2) + \delta)}
$$
infimum taken over the family of $(\epsilon,n)$-spanning sets $F \subset \Omega'$ and thus
$$
Q(f|\Omega',\phi,\epsilon) = \limsup_{n \to +\infty}\dfrac{\log{Q_n(f,\phi,\epsilon)}}{n} < P(f|\Omega,\phi,\epsilon/2) + \delta.
$$
Therefore, as $P(f|\Omega) = \sup_{\epsilon > 0}Q(f|\Omega,\phi,\epsilon) = \sup_{\epsilon > 0}P(f|\Omega,\phi,\epsilon)$ we conclude that
$$
P(f|\Omega',\phi) < P(f|\Omega,\phi) + \delta.
$$
By symmetry on $\Omega$ and $\Omega'$ we conclude that $|P(f|\Omega',\phi) - P(f|\Omega,\phi)| < \delta$ for $\dist_{H}(\Omega,\Omega') < \eta$. 

\end{proof}
 
\begin{proof}[Proof of Proposition \ref{prop.1}]
Clearly,
$$
P(f|\Lambda, \phi) \geq \sup_{\Omega \in \mathcal{H}}P(f|\Omega,\phi)
$$
since $P(f|\Lambda, \phi) \geq P(f|\Omega,\phi)$ for every compact $f$-invariant $\Omega \subset \Lambda$. On the other hand, by hypothesis and results in 
\cite{luzzatto.sanchez} there exists an exhaustion of $\Lambda$ by a sequence $\Lambda_n \subset \Lambda$ of basic sets such that
$\dist_H(\Lambda_n,\Lambda) \to 0^+$. Therefore, by the continuity of the pressure $P(f|\Omega,\phi)$ respect to $\Omega$,
$$
P(f|\Lambda_n, \phi) \to P(f|\Lambda, \phi),
$$
thus proving that
$$
P(f|\Lambda, \phi) = \sup_{\Omega \in \mathcal{H}}P(f|\Omega,\phi),
$$
where $\mathcal{H}$ is the family of basic sets in $\Lambda$.
\end{proof}

\begin{proof}[Proof of Theorem B]

Let $\bigwedge^{m}TM$ be the (one dimensional) vector bundle of volume forms over $M$, and $\bigwedge^{m}Df : \bigwedge^{m}TM \to \bigwedge^{m}TM$ 
the fiber map induced by the derivative $Df : TM \to TM$. Then, by Oseledec theorem and Kingman's subadditive ergodic theorem
$$
\Phi^u(x) = \lim_{n \to +\infty}-\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}(x)| = \sup_{n > 0}-\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}(x)| = -\sum_{\chi_i(x) > 0}\chi_i(x)\dim{E_i}(x)
$$
$\mu$-a.e. for every $f$-invariant Borel probability $\mu$.

Therefore,
$$
h(\mu) - \chi^+(\mu) = \sup_{n > 0}\left\{h(\mu) - \int\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}(x)|d\mu(x)\right\},
$$
where
$$
\chi^+(\mu) = \int\sum_{\chi_i(x) > 0}\chi_i(x)\dim{E_i}(x)d\mu(x).
$$
Let $\phi^u = -\log\text{Jac}(Df | E^u)$, the induced volume deformation along unstable manifolds. Then by Theorem C
\begin{equation}\label{eq.1}
\sup_{\mu \in \mathcal{M}_f(\Lambda)}\{h(\mu) - \chi^+(\mu)\} = \sup_{\Omega \in \mathcal{H}}P(f|\Omega,\Phi^u) 
\end{equation}
Indeed,
\begin{eqnarray*}
\sup_{\mu \in \mathcal{M}_f(\Lambda)}\{h(\mu) - \chi^+(\mu)\} & = & \sup_{\mu \in \mathcal{M}_f(\Lambda)}\sup_{n > 0}\left\{h(\mu) - \int\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}(x)|d\mu(x)\right\}\\
                                                              & = & \sup_{n > 0}\sup_{\mu \in \mathcal{M}_f(\Lambda)}\left\{h(\mu) - \int\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}(x)|d\mu(x)\right\}\\
                                                              & = & \sup_{n > 0}P\left(f|\Lambda,-\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}|\right)\\
                                                              & = & \sup_{n > 0}\sup_{\Omega \in \mathcal{H}}P\left(f|\Omega,-\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}|\right) \ \text{(by proposition \ref{prop.1})}\\
                                                              & = & \sup_{\Omega \in \mathcal{H}}\sup_{n > 0}P\left(f|\Omega,-\dfrac{1}{n}\log|\bigwedge^{m}Df^{n}|\right)\\
                                                              & = & \sup_{\Omega \in \mathcal{H}}P(f|\Omega,\Phi^u).
\end{eqnarray*}
On the other hand
$$
\lim_{n \to +\infty}\dfrac{1}{n}\phi^u(f^n(x)) = \Phi^u(x), \quad\quad\mu-a.e.\quad\forall \ \mu \in {\mathcal M}_f(\Omega),
$$
by Oseledec's theorem. Then, by Birkhoff's theorem
$$
\int\phi^u(x){d\mu(x)}  = \int\Phi^u(x)d\mu(x) \quad\quad\forall \ \mu \in {\mathcal M}_f(\Omega),
$$
since $\phi^u| \Omega$ is $\mu$-summable for every $f$-invariant Borel probability $\mu \in \mathcal{M}_f(\Omega)$.
Therefore, for every basic set $\Omega \subset \Lambda$,
$$
P(f|\Omega,\Phi^u) = P(f|\Omega,\phi^u) = \rho(U_{\Omega}),
$$
by the thermodynamics of Axiom A systems. See \cite[Proposition 4.8 (a)]{bowen}. This and equation (\ref{eq.1}) proves (\ref{eckmann.ruelle.2}) and, 
therefore, Theorem B.
\end{proof}

\section{When is regular a nonuniformly hyperbolic set?}

As we mentioned before Baladi et al. example shows that to be the support of nonatomic hyperbolic measure is not a sufficient condition for 
$\Lambda$ to be regular. Indeed let $U_0$ be the isolating neighborhood for $\Lambda_0$. Then $U_0 \subsetneq U$ is contained properly into $U$, the isolating neighborhood 
of $\Lambda$ and as long as $\Lambda_1$ is made up finitely many hyperbolic fixed points into an array of saddle connections and 
$\Lambda = closure \ (\Lambda_0)$ there is no chance to get a regular exhaustion of $\Lambda$.

We would like to conclude this note by proposing a sufficient condition for $\Lambda$ to be a regular nonuniformly hyperbolic set. But first we need 
some definitions.
\begin{definition}
Let $\Omega \subset M$ be a compact subset. We say that $\Omega$ has a \emph{hyperbolic product structure} if there exists two continuous laminations 
$\mathcal{F}^s$ and $\mathcal{F}^u$ of disks such that
\begin{enumerate}
\item the disks $\mathcal{F}^s(x)$ and $\mathcal{F}^u(x)$ have a size uniformly bounded between positive constants $\alpha < \beta$; 
\item $d(f^n(y),f^n(y')) \to 0^+$ (resp. $d(f^{-n}(z),f^{-n}(z')) \to 0^+$) exponentially fast for every $y,y' \in \mathcal{F}^s(x)$ (resp. $z,z' \in \mathcal{F}^u(x)$), for every $x \in \Omega$;
\item $\mathcal{F}^s$ and $\mathcal{F}^u$ are invariant: if $x \in \Omega$ and $f^n(x) \in \Omega$ then $f^n\mathcal{F}^s(x) \subset \mathcal{F}^s(f^n(x))$ and $f^{n}\mathcal{F}^u(x) \supset \mathcal{F}^u(f^n(x))$;
\item for every $x,y \in \Omega$ the disk $\mathcal{F}^s(x)$ intersects transversally $\mathcal{F}^u(y)$ at a unique point with an angle uniformly bounded from below by some constant $\gamma > 0$;
\item $\Omega = \bigcup\mathcal{F}^s \cap \bigcup\mathcal{F}^u$. 
\end{enumerate}
\end{definition}
See \cite[Definition 1]{young.1998}.

\begin{definition}
Let $\mu$ be a hyperbolic measure and $\Lambda = supp \ \mu$. We say that $\mu$ has a \emph{local hyperbolic product structure} if there exists Borel 
positive functions $\alpha,\beta, \gamma: \Lambda \to (0,+\infty)$ such that for every $x \in \Lambda$ there exists an open neighborhood $U_x$ and a 
compact set $\Omega \subset U_x$ with $\mu(\Omega) > 0$ having a hyperbolic product structure with constants $\alpha(x),\beta(x), \gamma(x)$.
\end{definition}
\ 
\\
{\bf Conjecture} \ {\em Let $\Lambda$ be an isolated nonuniformly hyperbolic set of a $C^{1+\alpha}$ diffeomorphism. Suppose that $\Lambda$ is 
the support a hyperbolic measure $\mu$ with positive entropy and local hyperbolic product structure. Then $\Lambda$ is regular. 
}
\ 
\\
\\
All the known systems satisfying Eckmann-Ruelle's conjecture have local hyperbolic product structure.

\end{document}